A and B of a sample space S, but assume that Pr[A]=0.2 and Pr[B]=0.6. Find Pr[A∪B] under each of the following conditions: (1) If A⊂B, then Pr[A∪B]= (2) If A∩B=∅, then Pr[A∪B]= (3) If A∩B′=∅, then Pr[A∪B]=
Answer #1

Given that, Pr[A] = 0.2 and Pr[B] = 0.6

We want to find, Pr[A∪B] under each of the following conditions.

1) If A⊂B, then Pr[A ∩ B] = Pr[A]

Pr[A∪B]

= Pr[A] + Pr[B] - Pr[A ∩ B]

= Pr[A] + Pr[B] - Pr[A]

= 0.2 + 0.6 - 0.2

= 0.6

=> Pr[A∪B] = 0.6

2) If A∩B=∅, then Pr[A∩B] = 0

Pr[A∪B] = Pr[A] + Pr[B] - Pr[A∩B] = 0.2 + 0.6 - 0

=> Pr[A∪B] = 0.8

3) If A∩B′=∅, then

Pr[A ∩ B'] = Pr[A] - Pr[A∩B]

=> 0 = Pr[A] - Pr[A ∩ B]

=> Pr[A ∩ B] = Pr[A]

Therefore,

Pr[A∪B]

= Pr[A] + Pr[B] - Pr[A ∩ B]

= Pr[A] + Pr[B] - Pr[A]

= 0.2 + 0.6 - 0.2

= 0.6

=> Pr[A∪B] = 0.6